### v2

@@ 3.2 A bound for $J(P)$ and $J_{\mathrm{w}}(P)$ using $\bar{\mathcal{R}}(C)$
In this section, we prove the following Theorem 3.5 that provides upper and lower bounds for $J(P)$ using the effective resistance of $\Phi(P^{*}P)$. Theorem 3.5 is a generalization of Theorem 3.1 in [21] for reversible $P$.
@@
To prove Theorem 3.5, in the same manner as the approach in [21], we define the weighted average effective resistance
@@
The similar statement appears in the proof of Lemma 5.8 of [21]. By changing $P^{2}$ to $P$ in [21], we obtain the lemma.
@@
Then we show the lower bound of $J_{\mathrm{w}}(P)$ in Theorem 3.5. When $PP^{*}=P^{*}P$, we can rewrite (14) as below:
@@ 3.3 A bound for $J(P)$ using the graph structure of network
Let $P$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then it holds that
@@
$J(P)\geq\frac{\pi_{\min}^{3}n}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)),\quad J_{\mathrm{w}}(P)\geq\frac{\pi_{\min}^{3}n^{2}}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
@@
By Theorem 3.5, it is enough for the proof to show the following lemma, which is a generalization of Lemma 5.9 in [21].
@@
Let $P$ be a consensus matrix with invariant measure $\bm{\pi}$ and $C_{P^{*}P}:=\Phi(P^{*}P)=nP^{\mathsf{T}}\varPi P$. Then
@@
Since the associated graph of the consensus matrix $P$ can be directed, we evaluate $\bar{\mathcal{R}}(C_{P^{*}P})$ using the maximum indegree $\delta_{\mathrm{in}}$ of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$, unlike previous research [21].
@@
In plain words, if the resistance of any edge increases, then the resistance between any pair of nodes also increases. The proof can be found in [10].
@@
Let $\mathcal{G}(P^{*}P)$ be the undirected graph associated with $P^{*}P$. Then, it holds that
@@
To compare the places of edges in $\mathcal{G}(P^{*}P)$ and $\mathcal{G}(P)$, we introduce the novel concepts of a back-and-forth path and its pivot node. A back-and-forth path between $u$ and $v$ is a path which consists of $\{u,w\},\{w,v\}\in\mathcal{G}(P)$ where $(w,u)$ and $(w,v)$ are edges in $\mathcal{G}_{\mathrm{dir}}(P)$. We refer to $w$ as the pivot node of the path (see Figure 2). When $\mathcal{G}_{\mathrm{dir}}(P)$ is symmetric, namely, $P_{ij}\neq 0$ if and only if $P_{ji}\neq 0$, back-and-forth paths reduce to paths of length $2$ in $\mathcal{G}(P)$. In such cases, including when $P$ is reversible, $\mathcal{G}(P^{*}P)$ has the same graph topology as $\mathcal{G}(P^{2})$. In general, the graph $\mathcal{G}(P^{h})$ is called $h$-fuzz, and previous research has shown the relationship between $\bar{\mathcal{R}}(\mathcal{G}(P))$ and $\bar{\mathcal{R}}(\mathcal{G}(P^{h}))$ [2]. However, for nonreversible $P$, it is worthwhile to compare $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$. Notice that the concept of a pivot node is close to that of a common in-neighbor, but differs slightly. Here, we emphasize that both the pivot node and the back-and-forth are concepts on the undirected graph, although they are defined via conditions on the directed graph.
@@
Let $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$ be the associated undirected graph with $P$ and $P^{*}P$, respectively. Then, it holds that
@@
For the left-hand inequality, we adapt the standard argument for fuzz graphs [2, Lem. 5.5.1]. Write $\mathcal{G}(P)=(V,\mathcal{E})$ and $\mathcal{G}(P^{*}P)=(V,\mathcal{E}^{*})$, and call edges in $\mathcal{E}^{*}\setminus\mathcal{E}$ “new edges” (see Figure 2). Replace each new edge $\{u,v\}$ by a series of two edges with an intermediate node $x_{uv}$, preserving the total resistance (e.g., split $r$ into $r/2+r/2$). This yields a graph $\bar{\mathcal{G}}=(\bar{V},\bar{\mathcal{E}})$ that is electrically equivalent to $\mathcal{G}(P^{*}P)$, so $\mathcal{R}(\bar{\mathcal{G}})=\mathcal{R}(\mathcal{G}(P^{*}P))$ (see Figure 3).
Next, for each new edge $\{u,v\}$, choose one back-and-forth path in $\mathcal{G}(P)$ with pivot $w$ and add a zero-resistance edge between $x_{uv}$ and $w$. Identifying these endpoints yields a graph $\mathcal{G}^{\prime}$ with $\mathcal{R}(\mathcal{G}^{\prime})\leq\mathcal{R}(\bar{\mathcal{G}})$ by Rayleigh’s monotonicity (see Figure 3) [^1] Intuitively, this operation can be understood as short-circuiting.. The graph $\mathcal{G}^{\prime}$ differs from $\mathcal{G}(P)$ only in the multiplicity of edges, so we compare the two by counting parallel edges. Note that $\mathcal{G}^{\prime}$ has one self-loop at each node, as does $\mathcal{G}(P)$. For $u,v\in V$ with $u\neq v$, the number of edges between $u$ and $v$ in $\mathcal{G}^{\prime}$ is bounded by the number of back-and-forth paths through $\{u,v\}$ in $\mathcal{G}(P)$. Let $\eta$ be an upper bound on the number of such paths. Since $k$ parallel unit-resistance edges are equivalent to a single edge of resistance $1/k$, Rayleigh’s monotonicity law yields $\mathcal{R}_{uv}\left(\mathcal{G}^{\prime}\right)\geq\frac{1}{\eta}\,\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)$. Combining this with the previous inequality, we obtain
@@
Substituting $P$ in Lemma 3.16 to $P^{*}$, then we obtain
@@
In addition, if $P^{*}P=PP^{*}$, we can change $\delta_{\mathrm{out}}$ in the leftmost term to $\delta_{\mathrm{in}}$ by Corollary 3.18. In this case, the lower bound for $\bar{\mathcal{R}}(C_{P^{*}P})$ can be obtained by $\frac{1}{n\pi_{\max}f(\delta_{\mathrm{in}})p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ where $f(\delta)=4\delta_{\mathrm{in}}^{2}+2\delta_{\mathrm{in}}-2$. Using Theorem 3.5, the proof is completed.
Theorem 3.11 follows immediately from Theorem 3.5 and Lemma 3.12.
@@ 4.1 Necessity of the assumption $P^{*}P=PP^{*}$
To understand the mechanism, note that, as $\varepsilon\to 0$, $P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially at a rate determined by the second largest eigenvalue of $P$. The eigenvalues of $P_{\varepsilon}$ are $1$ and $\alpha_{\pm}:=-\left(\frac{1}{4}-\varepsilon\right)\pm\frac{\mathrm{i}}{2}\sqrt{\frac{7}{4}-2\varepsilon}$. The eigenvalues of $P_{\varepsilon}^{*}$ coincide with those of $P_{\varepsilon}$ by similarity. Therefore, $(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially to some real matrix at a rate close to $1/2$. On the other hand, $P^{*}P$ has the second eigenvalue $1-O(\varepsilon)$. This means that if $\varepsilon$ is small enough, the convergence of $(P^{*}P)^{t}$ is very slow, and its trace can be much larger than $\operatorname{tr}(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}$. Therefore, the right-hand side of (16) becomes much larger than the leftmost hand, and Theorem 3.5 does not hold.
However, in Figure 4, it seems that the lower bound of Theorem 3.11 holds for small $\varepsilon$, although there are no theoretical guarantees in our paper. This is likely due to the rough evaluation of Lemma 3.12. When $\varepsilon$ decreases, the average effective resistance $\bar{\mathcal{R}}(C_{P^{*}P})$ increases linearly to $1/\varepsilon$, so the evaluation of Lemma 3.12 becomes rough. Taking into account this behavior, it may be possible to find sharper or more general lower bounds for $J(P)$ in future research.
@@ 4.2 Applications to Cayley graphs
An important application where the property $P^{*}P=PP^{*}$ holds is that of Cayley consensus matrices. Given a finite abelian group $(V,+)$ and a subset $S\subseteq V$, the Cayley graph $\mathcal{G}$ is defined as a directed graph with the vertex set $V$ and the edge set $E=\{(u,v)\mid u-v\in S\}$. The associated Cayley matrix $P\in\mathbb{R}^{V\times V}$ on this graph is defined by $P_{uv}=g(u-v)$ for some function $g$, called a generator, where rows and columns are indexed by the elements of $V$. If a consensus matrix $P$ is a Cayley matrix, then $P^{*}P=PP^{*}$ holds, since Cayley matrices are normal (see [5, Section III]) and Proposition 3.3 applies.
@@
Higher-dimensional tori are treated similarly using $\mathbb{Z}_{n}^{d}$. If we define $S$ as the neighbor set (for example, $\{-1,0,1\}^{d}$) and the value of $g(\bm{h})$ for each $\bm{h}\in S$ so that the sum of them is $1$, then we can consider a consensus algorithm on a $d$ dimensional grid where each agent communicates only with the neighbor agents. The example of a two dimensional torus is shown in Figure 6. In this case, we set $S=\{-1,0,1\}^{2}$ and $g(\pm 1,\pm 1)=0$. Grids with boundaries can be reduced to the torus case via an appropriate reflection argument [5].
@@
Proposition 4.1 (Proposition 1 in [5], modified).
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$, associated with a fixed generator $g$. Then there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
@@
Since $J(P)$ is already understood, we next relate $J(P)$ to our upper and lower bounds involving $\bar{\mathcal{R}}(\mathcal{G}(P))$.
@@
Let $P$ be a normal consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then it holds that
@@
Because a normal stochastic matrix is doubly stochastic, $\bm{\pi}=\frac{1}{n}\bm{1}$ and $\pi_{\min}=\pi_{\max}=\frac{1}{n}$. In addition, by the normality of $P$, $P^{*}=P$ holds and we have $P^{*}P=PP^{*}$. Applying them to Theorem 3.11, we complete the proof.
Corollary 4.2 establishes a direct relationship between the LQ cost $J(P)$ and the average effective resistance $\bar{\mathcal{R}}(\mathcal{G}(P))$ for normal consensus matrices. To further clarify the behavior of $\bar{\mathcal{R}}(\mathcal{G}(P))$ in concrete settings, we next specialize to Cayley graphs on $\mathbb{Z}_{n}^{d}$ and derive its asymptotic scaling.
Corollary 4.4.
Let $\mathcal{G}_{n}$ be the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$. Then there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
  • if $d=1$:
    $C_{d}n^{d}\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}n^{d}$
  • if $d=2$:
    $C_{d}\log(n^{d})\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}\log(n^{d})$
  • if $d\geq 3$:
    $C_{d}\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}$
Proof 4.5.
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on $\mathcal{G}_{n}$, associated with a fixed generator $g$. Then, $p_{\min},p_{\max}$ and an upper bound of $f(\delta_{\mathrm{in}})$ is determined by $g$. By Corollary 4.2, the upper and lower bounds of $J(P_{n})$ is linear to $\bar{\mathcal{R}}(\mathcal{G}_{n})$ under fixed $g$ and $d$. Therefore, substituting $J(P_{n})$ by $\bar{\mathcal{R}}(\mathcal{G}_{n})$ in Proposition 4.1, the proof is completed.
We compare $J(P)$ and $\bar{\mathcal{R}}(\mathcal{G}(P))$ as follows:
@@
  • Generate the Cayley graph $\mathcal{G}$ and the Cayley matrix $P\in^{n^{d}\times n^{d}}$ by the generator $g$. In case 1, $P$ is randomly determined, so we generate $20$ instances of $g$ for each $d,n$.
@@
Figure 9: The values of $J(P)$ in $d=2,3$, where $g$ is generated by the case 2. The left and right panels show the theoretical lower and upper bounds in Corollary 4.2.
The result is shown in Figure 7, 8, and 9. In these figures, $J(P)$ appears to increase logarithmically in $d=2$, and does not change much in $d=3$. In addition, the change of $\bar{\mathcal{R}}(\mathcal{G})$ seems to be consistent with Corollary 4.4. These results also imply that the bound we show in the paper is not sharp in some cases. In fact, the ratio of the lower and upper bounds in the case 1 of $d=3$ is $f(26)\times 0.1^{2}/0.01^{2}=275400$. However, in the case 2 of $d=2$, all elements are the same value, so the ratio is $f(2)=18$. Thus, our bounds are most informative when the in-degree is small and the generator weights are nearly uniform.

### v3

@@ 3.2 A bound for $J(P)$ and $J_{\mathrm{w}}(P)$ using $\bar{\mathcal{R}}(C)$
In this section, we prove the following Theorem 3.5 that provides upper and lower bounds for $J(P)$ using the effective resistance of $\Phi(P^{*}P)$. Theorem 3.5 is a generalization of Theorem 3.1 in [mainp] for reversible $P$.
@@
To prove Theorem 3.5, in the same manner as the approach in [mainp], we define the weighted average effective resistance
@@
The similar statement appears in the proof of Lemma 5.8 of [mainp]. By changing $P^{2}$ to $P$ in [mainp], we obtain the lemma.
@@
Then, we show the lower bound of $J_{\mathrm{w}}(P)$ in Theorem 3.5. When $PP^{*}=P^{*}P$, we can rewrite (14) as below:
@@ 3.3 A bound for $J(P)$ using the graph structure of network
Let $P$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then,
@@
$J(P)\geq\frac{\pi_{\min}^{3}n}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)),\quad J_{\mathrm{w}}(P)\geq\frac{\pi_{\min}^{3}n^{2}}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}}\bar{\mathcal{R}}(\mathcal{G}(P)),$
@@
By Theorem 3.5, it is enough for the proof to show the following lemma, which is a generalization of Lemma 5.9 in [mainp].
@@
Let $P$ be a consensus matrix with invariant measure $\bm{\pi}$ and $C_{P^{*}P}:=\Phi(P^{*}P)=nP^{\mathsf{T}}\varPi P$. Then,
@@
Since the associated graph of the consensus matrix $P$ can be directed, we evaluate $\bar{\mathcal{R}}(C_{P^{*}P})$ using the maximum indegree $\delta_{\mathrm{in}}$ of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$, unlike previous research [mainp].
@@
In plain words, if the resistance of any edge increases, then the resistance between any pair of nodes also increases. The proof can be found in [randomwalks].
@@
Let $\mathcal{G}(P^{*}P)$ be the undirected graph associated with $P^{*}P$. Then,
@@
To compare the places of edges in $\mathcal{G}(P^{*}P)$ and $\mathcal{G}(P)$, we introduce the novel concepts of a back-and-forth path and its pivot node. A back-and-forth path between $u$ and $v$ is a path which consists of $\{u,w\},\{w,v\}\in\mathcal{G}(P)$ where $(w,u)$ and $(w,v)$ are edges in $\mathcal{G}_{\mathrm{dir}}(P)$. We refer to $w$ as the pivot node of the path (see Figure 2). When $\mathcal{G}_{\mathrm{dir}}(P)$ is symmetric, namely, $P_{ij}\neq 0$ if and only if $P_{ji}\neq 0$, back-and-forth paths reduce to paths of length $2$ in $\mathcal{G}(P)$. In such cases, including when $P$ is reversible, $\mathcal{G}(P^{*}P)$ has the same graph topology as $\mathcal{G}(P^{2})$. In general, the graph $\mathcal{G}(P^{h})$ is called $h$-fuzz, and previous research has shown the relationship between $\bar{\mathcal{R}}(\mathcal{G}(P))$ and $\bar{\mathcal{R}}(\mathcal{G}(P^{h}))$ [hfuzz]. However, for nonreversible $P$, it is worthwhile to compare $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$. Notice that the concept of a pivot node is close to that of a common in-neighbor, but differs slightly. Here, we emphasize that both the pivot node and the back-and-forth are concepts on the undirected graph, although they are defined via conditions on the directed graph.
@@
Let $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$ be the associated undirected graph with $P$ and $P^{*}P$, respectively. Then,
@@
For the left-hand inequality, we adapt the standard argument for fuzz graphs [hfuzz, Lem. 5.5.1]. We write $\mathcal{G}(P)=(V,\mathcal{E})$ and $\mathcal{G}(P^{*}P)=(V,\mathcal{E}^{*})$, and call the edges in $\mathcal{E}^{*}\setminus\mathcal{E}$ “new edges” (see Figure 2). Consider replacing each new edge in $\mathcal{G}(P^{*}P)$ with a series of two edges and an intermediate node, resulting in a modified graph $\bar{\mathcal{G}}=(\bar{V},\bar{\mathcal{E}})$. In this transformation, the resistances of the new edges are doubled. By Rayleigh’s monotonicity law, the new resistance of every pair of two nodes in $V$ cannot exceed twice the original value, that is, $R_{uv}(\bar{\mathcal{G}})\leq 2R_{uv}(\mathcal{G}(P^{*}P))$ for every pair of nodes $u,v\in V$.
Next, focus on an intermediate node $x_{uv}$ added when replacing a new edge $\{u,v\}$. For every new edge $\{u,v\}$ in $\mathcal{G}(P^{*}P)$, there exists a back-and-forth path in $\mathcal{G}(P)$. Let $w$ be its pivot node. We then add a zero-resistance edge between $x_{uv}$ and $w$, and regard $x_{uv}$ and $w$ as an identical node. After this operation, the resulting graph $\mathcal{G}^{\prime}$ satisfies $\mathcal{R}_{uv}(\mathcal{G}^{\prime})\leq\mathcal{R}_{uv}(\bar{\mathcal{G}})$ for every pair of $u,v\in V$ by Rayleigh’s monotonicity (see Figure 3) [^1] Intuitively, this operation can be understood as short-circuiting.. The graph $\mathcal{G}^{\prime}$ differs from $\mathcal{G}(P)$ only in the multiplicity of edges, so we compare the two by counting parallel edges. Note that $\mathcal{G}^{\prime}$ has one self-loop at each node, as does $\mathcal{G}(P)$. For $u,v\in V$ with $u\neq v$, the number of edges between $u$ and $v$ in $\mathcal{G}^{\prime}$ is bounded by the number of back-and-forth paths through $\{u,v\}$ in $\mathcal{G}(P)$ plus $1$, the original edge. Let $\eta$ be an upper bound on the number of such paths plus the original edge. Since $k$ parallel unit-resistance edges are equivalent to a single edge of resistance $1/k$, Rayleigh’s monotonicity law yields $\mathcal{R}_{uv}\left(\mathcal{G}^{\prime}\right)\geq\frac{1}{\eta}\,\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)$. Combining this with the previous inequality, we obtain
@@
Substituting $P$ in Lemma 3.16 with $P^{*}$, then we obtain
@@
In addition, if $P^{*}P=PP^{*}$, we can change $\delta_{\mathrm{out}}$ in the leftmost term to $\delta_{\mathrm{in}}$ by Corollary 3.18. In this case, the lower bound for $\bar{\mathcal{R}}(C_{P^{*}P})$ can be obtained by $\frac{1}{n\pi_{\max}f(\delta_{\mathrm{in}})p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ where $f(\delta)=4\delta^{2}+2\delta-2$. Using Theorem 3.5, the proof is completed.
Proof 3.21 (Proof of Theorem 3.11).
Follows immediately from Theorem 3.5 and Lemma 3.12.
In the end of the section, we mention a special case in which our theorem can be written more simply.
Corollary 3.22.
Let $P$ be a normal consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then,
Proof 3.23.
Because a normal stochastic matrix is doubly stochastic, $\bm{\pi}=\frac{1}{n}\bm{1}$ and $\pi_{\min}=\pi_{\max}=\frac{1}{n}$. In addition, by Proposition 3.3, we have $P^{*}P=PP^{*}$. Using Theorem 3.11, we complete the proof.
@@ 4.1 Necessity of the assumption $P^{*}P=PP^{*}$
To understand the mechanism, note that, as $\varepsilon\to 0$, $P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially at a rate determined by the second largest eigenvalue of $P$. The eigenvalues of $P_{\varepsilon}$ are $1$ and $\alpha_{\pm}:=-\left(\frac{1}{4}-\varepsilon\right)\pm\frac{\mathrm{i}}{2}\sqrt{\frac{7}{4}-2\varepsilon}$. The eigenvalues of $P_{\varepsilon}^{*}$ coincide with those of $P_{\varepsilon}$ by similarity. Therefore, $(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially to some real matrix at a rate close to $1/2$. On the other hand, $P_{\varepsilon}^{*}P_{\varepsilon}$ has the second eigenvalue $1-O(\varepsilon)$. This means that if $\varepsilon$ is small enough, the convergence of $(P_{\varepsilon}^{*}P_{\varepsilon})^{t}$ is very slow, and $\operatorname{tr}(P_{\varepsilon}^{*}P_{\varepsilon})^{t}$ can be much larger than $\operatorname{tr}(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}$. Therefore, the right-hand side of (16) becomes much larger than the leftmost hand, and Theorem 3.5 does not hold.
However, in Figure 4, it seems that the lower bound of Theorem 3.11 holds for small $\varepsilon$, although there are no theoretical guarantees in our paper. This is likely due to the rough evaluation of Lemma 3.12. When $\varepsilon$ decreases, the average effective resistance $\bar{\mathcal{R}}(C_{P^{*}P})$ increases linearly to $1/\varepsilon$, making the evaluation in Lemma 3.12 increasingly loose. Taking into account this behavior, it may be possible to find sharper or more general lower bounds for $J(P)$ in future research.
@@ 4.2 Applications to Cayley graphs
An important application where the property $P^{*}P=PP^{*}$ holds is that of Cayley consensus matrices. Given a finite abelian group $(V,+)$ and a subset $S\subseteq V$, the Cayley graph $\mathcal{G}$ is defined as a directed graph with the vertex set $V$ and the edge set $E=\{(u,v)\mid u-v\in S\}$. The associated Cayley matrix $P\in\mathbb{R}^{V\times V}$ on this graph is defined by $P_{uv}=g(u-v)$ for some function $g$, called a generator, where rows and columns are indexed by the elements of $V$. If a consensus matrix $P$ is a Cayley matrix, then $P^{*}P=PP^{*}$ holds, since Cayley matrices are normal (see [cayley, Section III]) and Proposition 3.3 applies.
@@
Higher-dimensional tori are treated similarly using $\mathbb{Z}_{n}^{d}$. If we define $S$ as the neighbor set (for example, $\{-1,0,1\}^{d}$) and the value of $g(\bm{h})$ for each $\bm{h}\in S$ so that the sum of them is $1$, then we can consider a consensus algorithm on a $d$ dimensional grid where each agent communicates only with the neighbor agents. The example of a two dimensional torus is shown in Figure 6. In this case, we set $S=\{-1,0,1\}^{2}$ and $g(\pm 1,\pm 1)=0$. Grids with boundaries can be reduced to the torus case via an appropriate reflection argument [cayley].
@@
Proposition 4.1 (Proposition 1 in [cayley], modified).
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$, associated with a fixed generator $g$. Then, there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
@@
Since $J(P)$ is already understood, we next relate $J(P)$ to our upper and lower bounds involving $\bar{\mathcal{R}}(\mathcal{G}(P))$. Corollary 3.22 establishes a direct relationship between the LQ cost $J(P)$ and the average effective resistance $\bar{\mathcal{R}}(\mathcal{G}(P))$ for normal consensus matrices. To further clarify the behavior of $\bar{\mathcal{R}}(\mathcal{G}(P))$ in concrete settings, we specialize to Cayley graphs on $\mathbb{Z}_{n}^{d}$ and derive its asymptotic scaling.
@@
Let $\mathcal{G}_{n}$ be the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$. Then, there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
  • if $d=1$:
    $C_{d}n^{d}\leq\bar{\mathcal{R}}(\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}})\leq C_{d}^{\prime}n^{d}$
  • if $d=2$:
    $C_{d}\log(n^{d})\leq\bar{\mathcal{R}}(\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}})\leq C_{d}^{\prime}\log(n^{d})$
  • if $d\geq 3$:
    $C_{d}\leq\bar{\mathcal{R}}(\mathcal{G}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}})\leq C_{d}^{\prime}$
@@
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on $\mathcal{G}_{n}$, associated with a fixed generator $g$. Then, $p_{\min},p_{\max}$ and an upper bound of $f(\delta_{\mathrm{in}})$ is determined by $g$. By Corollary 3.22, the upper and lower bounds of $J(P_{n})$ depend linearly on $\bar{\mathcal{R}}(\mathcal{G}_{n})$ under fixed $g$ and $d$. Therefore, substituting $J(P_{n})$ with $\bar{\mathcal{R}}(\mathcal{G}_{n})$ in Proposition 4.1, the proof is completed.
Thus, not only $J(P)$ but also the bounds of $J(P)$ in Corollary 3.22 grow logarithmically in $d=2$ and keep constant in $d=3$. To check these behaviors, we calculate $J(P)$ and the lower and upper bounds in Corollary 3.22 as follows:
@@
  • Generate the Cayley graph $\mathcal{G}$ and the Cayley matrix $P\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}}^{n^{d}\times n^{d}}$ by the generator $g$. In case 1, $P$ is randomly determined, so we generate $20$ instances of $g$ for each $d,n$.
@@
Figure 9: The values of $J(P)$ in $d=2$ (left panel) and $d=3$ (right panel), where $g$ is generated by the case 2. The left and right panels show the theoretical lower and upper bounds in Corollary 3.22.
The result is shown in Figure 7, 8, and 9. In these figures, $J(P)$ appears to increase logarithmically in $d=2$, and does not change much in $d=3$. In addition, the change of $\bar{\mathcal{R}}(\mathcal{G})$ seems to be consistent with Corollary 4.2. These results also imply that the bound we show in the paper is not sharp in some cases. In fact, the ratio of the lower bound to the upper bound in the case 1 of $d=3$ is $f(26)\times 0.1^{2}/0.01^{2}=275400$. However, in the case 2 of $d=2$, all elements are the same value, so the ratio is $f(2)=18$. Thus, our bounds are most informative when the in-degree is small and the generator weights are nearly uniform.